### v3

@@ 4.3 Ablation analysis
We further compare our benchmark against extreme multi-class classification settings. The prompt template is shown in Figure 11 in the Appendix, and we select the best-performing model from each category. The evaluation uses triplet-level ($\{\texttt{Tag},\texttt{Fact},\texttt{Type}\}$) Precision, Recall, and F1.
@@
Table 8: Performance comparison between with/without our evaluation framework on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark dataset.
@@
As shown in Table 8, treating XBRL tagging as single-step extreme classification fails in practice. The model must choose directly from thousands of taxonomy concepts without first identifying valid numerical entities, forcing it to rely on internal priors rather than grounded evidence. All models collapse to zero, illustrating that this formulation is not a meaningful way to evaluate LLMs on XBRL tagging. In contrast, our two-stage framework mirrors how tagging is actually performed. FinNI first detects and types candidate numerical facts, and FinCL then retrieves a small set of plausible tags and requires the model to select among them using context. Under this formulation, models achieve non-trivial performance, revealing real differences in capability that are completely hidden under extreme classification.
4.4 Error propagation in 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: pipeline
Table 9 compares accuracy across three settings: the full pipeline, FinCL with gold entities, and rerank-only with guaranteed gold candidates. The pattern reveals clear cascading errors. End-to-end performance is consistently the lowest because mistakes from extraction and retrieval accumulate in the linking stage.
When gold entities are supplied, accuracy improves substantially, showing that entity identification is a major bottleneck. Performance increases again in the rerank-only setting, indicating that narrowing the label space reduces difficulty. However, accuracy still remains far from solved, suggesting that fine-grained US-GAAP disambiguation is intrinsically challenging even when retrieval is perfect. In addition, we provide representative error cases in Appendix I, illustrating how these challenges arise in practice.
Table 9: Accuracy comparison across the full pipeline, FinCL subtask, and rerank-only settings.
@@ 5 Conclusion
Limitations
In this work, we propose 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and conduct a comprehensive analysis of LLM performance on the task. However, several limitations remain. First, we have not yet evaluated the latest models, such as GPT-5.2, gpt-oss, and Qwen3-235B. We are currently conducting these experiments and will report the results in future work. Second, we collected only a small dataset for fine-tuning PLMs and did not construct task-specific training data for LLMs. Future work may explore building LLM-based XBRL tagging agents through targeted fine-tuning to enhance performance further.
Potential Risks
Potential Positive Societal Impacts.
0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides the first structure-aware benchmark for evaluating large language models on realistic XBRL tagging. By reframing tagging as a transparent, two-stage Extract–Align process, this benchmark:
  • Enables more rigorous and reproducible evaluation of model performance on financial fact extraction and taxonomy alignment.
  • Supports research on safer automation of financial workflows by exposing failure modes that arise in long-tailed concept spaces and complex tabular contexts.
  • Lowers the barrier for auditing, interpreting, and improving financial AI systems by releasing open datasets, evaluation protocols, and diagnostic analyses.
By emphasizing open evaluation rather than proprietary pipelines, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: contributes to more transparent and accountable financial AI research.
Potential Negative Societal Impacts.
Despite these benefits, several risks remain:
  • Misuse and Overreliance. High benchmark performance may be mistakenly interpreted as readiness for real-world XBRL tagging, regulatory reporting, or financial decision-making, even though the benchmark does not replace human review or auditing.
  • Bias and Coverage Limitations. The dataset is drawn from U.S. SEC filings and US-GAAP concepts. As a result, reporting practices, sector distributions, and taxonomy frequency patterns may introduce biases and may not generalize to other jurisdictions or accounting standards.
  • Benchmark Narrowness. 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: evaluates extraction and alignment only. Downstream tasks such as aggregation, reconciliation, audit validation, and materiality assessment remain outside the benchmark and may still fail in deployed systems.
Ongoing iteration, external feedback, and careful documentation are necessary to mitigate these concerns.
Ethical Considerations and Usage Disclaimer
The authors assume responsibility for the design and dissemination of 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. All data used in the benchmark are derived from publicly available SEC filings and contain no intentionally included personal or sensitive information. Dataset construction followed community ethical guidelines for research involving public financial disclosures.
Unless otherwise specified, the datasets and accompanying evaluation resources are released for research and educational purposes. They should not be used as a substitute for professional accounting judgment, financial reporting systems, or regulatory compliance procedures.
The benchmark is designed to evaluate model behavior under controlled conditions. Performance on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: does not imply suitability for production deployment, financial auditing, or automated investment or compliance tasks. Any real-world use must include human oversight, documentation of limitations, and appropriate legal and regulatory review.
We used AI-assisted writing tools to help refine the wording and improve the readability of the manuscript. All scientific claims, dataset design choices, and experimental analyses were developed and verified by the authors.
While reasonable care has been taken to ensure accuracy, the authors provide no warranty regarding completeness or suitability for any particular purpose and disclaim liability for any damages arising from use of the benchmark or associated materials.
Licensing and Terms of Use
We release all benchmark data, annotations, and accompanying resources under the CC-BY 4.0 license, which permits redistribution and adaptation with attribution. All source documents originate from publicly available SEC filings, and we strictly comply with the original terms governing their public distribution. The US-GAAP taxonomy is used in accordance with its publicly documented terms. Our released benchmark does not contain proprietary material beyond what is already publicly accessible, and any derivative use must continue to acknowledge both the original providers and our benchmark.
@@ Appendix A Significance analysis
Figure 4: Visualization of significance analysis across models. (a) shows the pairwise-bootstrap significance matrix, while (b) summarizes how many significant differences each model has relative to others.
@@ Appendix B Literature Review
B.1 XBRL Tagging benchmark
@@
FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging loukas2022finer. The initial experiments showed that standard BERT underperforms due to subword fragmentation, then the authors introduced pseudo-token strategies replacing numerals with [NUM] or [SHAPE] tokens to stabilize label assignment across fragmented numeric spans. These strategies, combined with domain-specific pretraining on SEC-BERT, significantly improved tagging performance, reaching 82.1 micro-F1 without the need for computationally expensive CRF layers. Their experiments demonstrated that subword-aware models with numeric-aware pseudo-tokens outperform both word-level BiLSTMs and vanilla BERT, particularly in numeric-heavy contexts, and avoid nonsensical label sequences. FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem sharma2023financial. They compared the FiNER sequence labeling approach with a two-step pipeline that first identifies numeric spans and then assigns labels using AttentionXML. While FiNER achieved stronger micro-F1 (75.84), reflecting better performance on frequent tags, AttentionXML outperformed FiNER in macro-F1 (47.54), highlighting its strength in predicting infrequent, tail-end labels. FNXL further evaluated both models under a Hits@k setting, confirming that label recommendations from the AttentionXML pipeline could substantially reduce manual effort and maintain high inter-annotator agreement. Together, these benchmarks reveal the need for context-aware reasoning and label-ranking mechanisms in realistic XBRL tagging scenarios.
B.2 XBRL Tagging Methods
@@
The previous studies also explored approaches to address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve performance. Saini et al. saini2021galaxc proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods. By integrating label metadata and transitive label correlations, GalaXC outperformed leading deep classifiers by up to 18% in micro-F1 on standard benchmarks and achieved 25% gains in warm-start scenarios where partial labels are available. Moreover, Wang et al. wang2023standardizing addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings. Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements, offering a low-cost and interpretable solution for downstream financial analysis. Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports han2024xbrlagent. The authors introduced two task types, domain taxonomy understanding and numeric reasoning, and found that base LLMs often hallucinated or misinterpreted financial content. To overcome these issues, XBRL-Agent incorporated retrieval-augmented generation (RAG) and symbolic calculators within an LLM-agent framework. The enhanced system achieved a 17% accuracy gain on domain query tasks and a 42% boost on numeric reasoning queries compared to base LLMs, validating the utility of modular tool augmentation. These improvements enabled reliable multi-step reasoning over complex disclosures such as debt instruments and derivative gains, which are difficult to capture using span-level classifiers. Collectively, these works broaden the methodological landscape of XBRL tagging from graph-based label propagation and embedding-based normalization to LLM-driven report analysis and point to a hybrid future where structured priors and reasoning tools jointly support accurate, scalable financial information extraction.
B.3 Financial Evaluation Benchmarks
@@
In parallel to XBRL-specific advances, the financial NLP community has developed comprehensive benchmarks to assess broader capabilities in information extraction, numerical reasoning, and document understanding. FiNER-ORD shah2023finer introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations. While not numerically focused like FiNER-139, it highlights the lexical diversity of financial discourse and establishes a strong baseline for testing pretrained and zero-shot LLMs in real-world financial NER scenarios. FinQA chen2021finqa pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports. ConvFinQA chen2022convfinqa extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries. TAT-QA zhu2021tat focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation. Pixiu xie2023pixiu introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction. BizBench koncel2023bizbench framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid. While these benchmarks highlight the growing ability of models to integrate structured and unstructured financial data, they overlook taxonomy-driven fact alignment and do not support the structured output formats required for XBRL tagging.
@@ Appendix C Data and Ticker Information
C.1 The Statistics of the Tickers
Figure 5, Figure 6, and Figure 7 illustrate the distribution of 142 tickers across industry sectors, market capitalization categories, and geographic regions, highlighting the diversity of our collected financial reports. Considering that practical XBRL tagging practices often vary by industry, company size, and legal jurisdiction, we curated a diverse set of companies covering all 11 major industry sectors. The sample maintains a balanced distribution across firm sizes, following market capitalization categories: micro-cap (<\$300M), small-cap (\$300M–\$2B), mid-cap (\$2B–\$10B), large-cap (\$10B–\$200B), and mega-cap (>\$200B). In addition, we incorporated firms from over 30 states and international jurisdictions to capture regional differences in reporting and tagging conventions. This diversity ensures that the benchmark reflects realistic heterogeneity observed in financial disclosures across sectors, scales, and regulatory environments.
Figure 5: Distribution of Tickers by Industry Sector.
Figure 6: Distribution of Tickers by Market Cap Size.
Figure 7: Distribution of Tickers by Geography.
C.2 Data Granularity Justification

### v4

@@ 4.3. Ablation analysis
We further compare our benchmark against extreme multi-class classification settings. We select the best-performing model from each category. The evaluation uses triplet-level ($\{\texttt{Tag},\texttt{Fact},\texttt{Type}\}$) Precision, Recall, and F1.
@@
Table 7. Performance comparison between with/without our evaluation framework on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark dataset.
@@
Table 7 shows that modeling XBRL tagging as single-step extreme classification fails in practice: models must select from thousands of taxonomy concepts without grounding in identified numerical entities, leading all models to collapse to zero. In contrast, our two-stage framework reflects real tagging workflows: FinNI first identifies and types numerical facts, and FinCL then selects from a retrieved candidate set using context. This formulation yields non-trivial performance and exposes meaningful differences among models that are entirely obscured under extreme classification.
4.4. Error propagation in 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: pipeline
Table 8 compares accuracy under three settings: the full pipeline, FinCL with gold entities, and rerank-only with gold candidates. Results show clear error cascading: end-to-end performance is lowest due to accumulated extraction and retrieval errors. Providing gold entities substantially improves accuracy, identifying entity extraction as a key bottleneck. Rerank-only settings further increase performance, yet accuracy remains low, indicating that fine-grained US-GAAP disambiguation is inherently challenging even with perfect retrieval.
Table 8. Accuracy comparison across the full pipeline, FinCL subtask, and rerank-only settings.
@@ Appendix A Significance analysis
Figure 4. Visualization of significance analysis across models. (a) shows the pairwise-bootstrap significance matrix, while (b) summarizes how many significant differences each model has relative to others.
@@ Appendix B Literature Review
B.1. XBRL Tagging benchmark
@@
FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging (Loukas et al., 2022). The initial experiments showed that standard BERT underperforms due to subword fragmentation, then the authors introduced pseudo-token strategies replacing numerals with [NUM] or [SHAPE] tokens to stabilize label assignment across fragmented numeric spans. These strategies, combined with domain-specific pretraining on SEC-BERT, significantly improved tagging performance, reaching 82.1 micro-F1 without the need for computationally expensive CRF layers. Their experiments demonstrated that subword-aware models with numeric-aware pseudo-tokens outperform both word-level BiLSTMs and vanilla BERT, particularly in numeric-heavy contexts, and avoid nonsensical label sequences. FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem (Sharma et al., 2023). They compared the FiNER sequence labeling approach with a two-step pipeline that first identifies numeric spans and then assigns labels using AttentionXML. While FiNER achieved stronger micro-F1 (75.84), reflecting better performance on frequent tags, AttentionXML outperformed FiNER in macro-F1 (47.54), highlighting its strength in predicting infrequent, tail-end labels. FNXL further evaluated both models under a Hits@k setting, confirming that label recommendations from the AttentionXML pipeline could substantially reduce manual effort and maintain high inter-annotator agreement. Together, these benchmarks reveal the need for context-aware reasoning and label-ranking mechanisms in realistic XBRL tagging scenarios.
B.2. XBRL Tagging Methods
@@
The previous studies also explored approaches to address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve performance. Saini et al. (Saini et al., 2021) proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods. By integrating label metadata and transitive label correlations, GalaXC outperformed leading deep classifiers by up to 18% in micro-F1 on standard benchmarks and achieved 25% gains in warm-start scenarios where partial labels are available. Moreover, Wang et al. (Wang, 2023) addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings. Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements, offering a low-cost and interpretable solution for downstream financial analysis. Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports (Han et al., 2024). The authors introduced two task types, domain taxonomy understanding and numeric reasoning, and found that base LLMs often hallucinated or misinterpreted financial content. To overcome these issues, XBRL-Agent incorporated retrieval-augmented generation (RAG) and symbolic calculators within an LLM-agent framework. The enhanced system achieved a 17% accuracy gain on domain query tasks and a 42% boost on numeric reasoning queries compared to base LLMs, validating the utility of modular tool augmentation. These improvements enabled reliable multi-step reasoning over complex disclosures such as debt instruments and derivative gains, which are difficult to capture using span-level classifiers. Collectively, these works broaden the methodological landscape of XBRL tagging from graph-based label propagation and embedding-based normalization to LLM-driven report analysis and point to a hybrid future where structured priors and reasoning tools jointly support accurate, scalable financial information extraction.
B.3. Financial Evaluation Benchmarks
@@
In parallel to XBRL-specific advances, the financial NLP community has developed comprehensive benchmarks to assess broader capabilities in information extraction, numerical reasoning, and document understanding. FiNER-ORD (Shah et al., 2023) introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations. While not numerically focused like FiNER-139, it highlights the lexical diversity of financial discourse and establishes a strong baseline for testing pretrained and zero-shot LLMs in real-world financial NER scenarios. FinQA (Chen et al., 2021) pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports. ConvFinQA (Chen et al., 2022) extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries. TAT-QA (Zhu et al., 2021) focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation. Pixiu (Xie et al., 2023) introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction. BizBench (Koncel-Kedziorski et al., 2023) framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid. While these benchmarks highlight the growing ability of models to integrate structured and unstructured financial data, they overlook taxonomy-driven fact alignment and do not support the structured output formats required for XBRL tagging.
@@ Appendix C Data and Ticker Information
C.1. The Statistics of the Tickers
Figure 5, Figure 6, and Figure 7 illustrate the distribution of 142 tickers across industry sectors, market capitalization categories, and geographic regions, highlighting the diversity of our collected financial reports. Considering that practical XBRL tagging practices often vary by industry, company size, and legal jurisdiction, we curated a diverse set of companies covering all 11 major industry sectors. The sample maintains a balanced distribution across firm sizes, following market capitalization categories: micro-cap (¡\$300M), small-cap (\$300M–\$2B), mid-cap (\$2B–\$10B), large-cap (\$10B–\$200B), and mega-cap (¿\$200B). In addition, we incorporated firms from over 30 states and international jurisdictions to capture regional differences in reporting and tagging conventions. This diversity ensures that the benchmark reflects realistic heterogeneity observed in financial disclosures across sectors, scales, and regulatory environments.
Figure 5. Distribution of Tickers by Industry Sector.
Figure 6. Distribution of Tickers by Market Cap Size.
Figure 7. Distribution of Tickers by Geography.
C.2. Data Granularity Justification